OpenStudy (anonymous):

solve 1/3x+9 - 2/x+3 = 2

OpenStudy (anonymous):

Is that\[\frac{ 1 }{ 3x + 9 } - \frac{ 2 }{ x + 3 } = 2\]? It's hard to tell with the way you wrote it just what's in the numerator and denominator.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ x + 3 }{ (3x + 9)(x + 3) } - \frac{ 2(3x + 9) }{ (x + 3)(3x + 9) } = 2\]You now have a common denominator for the 2 terms on the left. Can you handle it from here or do you want more help?

OpenStudy (anonymous):

i think i got it. dont i cancel them out then it will be -1/ (3x+9)(x+3) = 2

OpenStudy (anonymous):

\[\frac{ (x + 3) - (6x + 18) }{ (3x + 9)(x + 3) } = 2\]\[\frac{ -5x - 15 }{ (3x + 9)(x + 3) } = 2\]\[\frac{ -5\cancel{(x + 3)} }{ (3x + 9)\cancel{(x + 3)} } = 2\]

OpenStudy (anonymous):

-5 = 2(3x + 9)

OpenStudy (anonymous):

then multiply both sides by the 3x+9 and then simplify?

OpenStudy (anonymous):

yes, you got it!

OpenStudy (phi):

it would be better to notice that 3x+9 is 3(x+3) your common denominator is 3(x+3) multiply the 2nd fraction by 3/3

OpenStudy (anonymous):

Good luck to you in all of your studies and thax for the recognition! @minimoo

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

uw!